# SolNet: Open-source deep learning models for photovoltaic power forecasting across the globe

### v2

@@ Abstract
Deep learning models have gained increasing prominence in recent years in the field of solar photovoltaic (PV) forecasting. One drawback of these models is that they require a lot of high-quality data to perform well. This is often infeasible in practice, due to poor measurement infrastructure in legacy systems and the rapid build-up of new solar systems across the world. This paper proposes SolNet: a novel, general-purpose, multivariate solar power forecaster, which addresses these challenges by using a two-step forecasting pipeline which incorporates transfer learning from abundant synthetic data generated from PVGIS, before fine-tuning on observational data. Using actual production data from hundreds of sites in the Netherlands, Australia and Belgium, we show that SolNet improves forecasting performance over data-scarce settings as well as baseline models. We find transfer learning benefits to be the strongest when only limited observational data is available. At the same time we provide several guidelines and considerations for transfer learning practitioners, as our results show that weather data, seasonal patterns, amount of synthetic data and possible mis-specification in source location, can have a major impact on the results. The SolNet models created in this way are applicable for any land-based solar photovoltaic system across the planet where simulated and observed data can be combined to obtain improved forecasting capabilities.
@@ 1.2 Contributions
Second, we do a thorough deep-dive of various transfer learning scenarios, providing guidance on when and how to use it successfully in solar power generation scenario’s. We look at three performance metrics for better understanding of the value of transfer: initial performance, learning improvement, and asymptotic performance. Figure 1 visualises this idea. Additionally, we look at (1) the importance of weather data, (2) the impact of the yearly seasonality on transfer learning out-performance, and (3) how mis-specification of the source domain can impact the performance of transfer learning.
Figure 1: Complete assessment of transfer performance. The initial performance measures performance of a model without any training in the domain of interest (the target domain). The learning performance indicates the performance gain we can expect over time as more data to train on becomes available. Finally, the asymptotic performance is the ’permanent’ performance differential between the models when sufficient data is available.
@@ 2.1 The forecasting exercise and baseline approach
For this paper we focused on the case of forecasting the next 24 hours, from 00:00 to 23:00 with no gap between inference time and the forecast period. This is a simple reference case to showcase the performance of SolNet. The main reason for choosing this option is that, when using weather forecast data (which we get from the Deutscher Wetterdienst’s Icosahedral Nonhydrostatic EU model (DWD ICON EU) via Open Climate Fix (Open Climate Fix, 2023)), we have to decide on a reference timing. Historical weather forecasts, when and if they are available, typically offer the option to choose from four different reference time slots: 00:00, 06:00, 12:00 and 18:00. We limited ourselves to the 00:00 data, as otherwise the dimensionality of the data would have become too large to work with.
@@ 2.2 LSTM as a forecast model
$h_{t-1}$ : is the hidden state at time $t-1$
$c_{t-1}$ : is the cell state at time $t-1$
@@ 2.3 Transfer learning
With the deep learning architecture in place, it is possible to explore the effect of transfer learning on solar forecasting in a principled manner. More concretely, the LSTM can be pre-trained on source data, after which it can be fine-tuned using target data to achieve potential improvements over the target model. Transfer learning, formally defined as improving the learning of a target predictive function on the learning task, $\mathcal{F}_{T}(\cdot)$ in the target domain, $\mathcal{D}_{T}$, using knowledge of $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$ (the source domain and the source task respectively) (Pan & Yang, 2010). While $\mathcal{D}_{T}\neq\mathcal{D}_{S}$ and $\mathcal{T}_{T}\neq\mathcal{T}_{S}$, we do assume that, the more similar the two are, the easier the transfer learning task will be. In our case, we keep the source domain as closely related to the target domain as possible by mimicking the location and the PV set-up of our target domain data, in our source domain, but we also provide results for what happens when this condition does not hold (see section 4.5). This is in addition to the inherent mis-alignment between simulated PVGIS data and observed data.
@@
    Finetuning. The source LSTM can be fine-tuned to the target domain. We keep the architecture intact, and run several additional training iterations using the (limited) target domain data. We can avoid model overfitting on the limited target data by lowering the learning rate, thereby making sure that the model only makes slight adjustments to the previously trained weights. Another popular method to fine-tune deep neural networks is by ”freezing” the parameters of certain units or entire layers (see e.g. Zhou et al. (2020b)). In this case, the optimiser can no longer update the parameters in these layers during the optimisation process, but is free to update the remaining parameters.
@@ 2.4 Model evaluation
Performance is measured by way of the root mean squared error (RMSE), in accordance with (Yang et al., 2020), since it penalizes large errors more heavily. This is important to correctly predict the timing and size of the daily peak in output. Another reason to use the RMSE is its ubiquity in solar forecasting literature, which makes it easier to compare across studies. In all cases, a full year of target data is kept aside for final evaluation. 80% of the remaining data is used for training and 20% is used to validate the model performance and pick the best performing iteration of the model.
The complete workflow of the experiments is presented in figure 2. The first column shows the workflow for a source trained model. The middle column shows the process for a target trained model, skipping the source step and training directly on the target domain instead of finetuning with a lower learning rate and frozen layers. The final column shows the naive seasonal model, which takes the values of the day before as the predictions, requiring no training.
@@ 3.1.1 Residential buildings in the Netherlands
For all 40 locations, the full year of 2020 is available as well as 2021 up until August. We also have a part of 2019 for several locations, but unfortunately our weather forecast data is limited to 2020 and 2021, so we cannot go back further for target locations if we want to compare univariate and multivariate performance. As mentioned, a full year of data is kept aside for evaluation of the final models. Which will thus consist of a part of 2021 and a part of 2020. This leaves little data for training and validation in the target domain, but this fits well with the core premise of the research that there is often only a (very) limited amount of data available in the target domain before forecasts need to be made.
@@ 3.2 Source domain
For this research we use data generated by PVGIS from 2005 to 2019, mimicking the parameters (latitude, longitude, tilt, azimuth, peak power) of each of the previously discussed solar PV systems (in case these parameters are available) in order to maximize the similarity between source and target domain. So, for example, an actual PV system in EnergyVille, Genk, Belgium with latitude 50.99 and longitude 5.54 a tilt of 33°, point directly south with a peak power of 2.5 kWp would get these exact parameters in the PVGIS tool to generate data from 2005 until 2019.
This data is subsequently used to train the source models we use for transfer learning. These models are then fine-tuned using the sparse observational target data. Even though PVGIS provides data for 2020, we do not use this year while creating source models as this can potentially cause data leakage. Data leakage refers to including knowledge about the future (or data which would realistically not be available) in a model. In this case, data leakage would occur if we were to use the year 2020 in the creation of our model and, simultaneously, evaluate the model on 2020. Similarly, for the Australian PV systems, for which we have data from 2011 to 2012, we limit the source data until 2010.
@@ 3.3 Weather data
To mitigate this issue, we treat the source and target domain differently. For training the source model, we utilize readily available observed historical weather data. For the target models, we only utilize actual forecast weather data. While this introduces a mis-alignment between source and target models, it does not cause any data leakage issues and also avoids placing large and potentially expensive forecast availability requirements on SolNet users, or anyone looking to use transfer-learning for their solar power forecasts.
@@ 3.4 Data pre-processing
[…]
@@ 4 Results
In this section, we present results of applying transfer learning to residential PV systems in the Netherlands and Australia, as well as the commercial PV systems in Belgium using univariate data. Next, using the Dutch data, we take a closer look at the impact of including weather covariates. Then we look at the importance of the time of the year when we start gathering target data on model performance. To do this, we change the training period to assess the seasonal impact. Finally, we also assess the impact of data mis-specification in the source domain, by increasing the (real) distance between the source and target domain.
4.1 Univariate forecasting of residential PV systems
The results for univariate forecasting of 40 residential PV systems in the Netherlands are summarised in figure 3(a) (i.e. without incorporating weather observations or forecasts in either the source or target domain). On the x-axis we plot the performance based on the months of data available in the target domain. ”0m” is the case with no target data (cf. 1. Likewise, ”1m” means that one month of target domain data is used for training/finetuning. On the y-axis we plot the scaled RMSE for each of these cases (i.e. the RMSE on the data which has been scaled using the min-max scaler). The thick lines are the overall median [^2] The median values are shown instead of mean values because not all datasets have the same observation period, which greatly impacts the mean estimate. values and the shaded areas represent the range of the individual cases. In other words, the shaded area represents the results of all 40 PV systems, from lowest to highest RMSE. The error of the baseline model (a seasonal naive model taking the values of the 24 hours of the previous day to be the new prediction), is a straight line, which does not improve over time, as it is independent of the amount of data available.
@@ 4.1 Univariate forecasting of residential PV systems
Figure 3: Forecast error as a function of target data availability and use of transfer learning for univariate models. Results indicate transfer learning models consistently outperform both naive baselines and target models (i.e. trained without transfer)
@@
From this plot, we can make four observations. First, the transfer models strictly outperform the baseline models. Second, the target model only starts to outperform the simple baseline once a sufficient amount of data (approx. 4 to 6 months of data depending on the test site) is available to train the model. This is very likely driven by the fact that this time period is sufficient for the model to have seen months with different weather conditions allowing it to generalize. Third, we can see that the asymptotic performance improvement is almost non-existent, with the RMSE staying at roughly 0.115. Fourth, the transfer model almost always outperforms the target model, until more than a half year of data is available.
@@
Figure 3(b) plots the corresponding results for Australia. Given the larger size of the dataset (299 samples), the shaded area shows the 0.05 to 0.95 percentiles for each month to filter out some of the noise in the outliers. In a similar fashion to the Dutch results, most of the value lies in the initial months. As before, it takes over 7 months for the target model to consistently outperform the naive model here. The learning improvement performance for the transfer model is again small. The median RMSE is 0.135 for the model with no fine-tuning (0m), improving by about 7.4% to 0.125 after 12 months. The target models outperform the transfer models in 8.6% of cases, mostly in the final four months.
4.2 Univariate forecasting of commercial PV systems
@@
Figure 3(c) plots the results for the two commercial buildings in Belgium, showing that results are in line with the residential results. As before, the transfer model outperforms the target model, with most of the performance difference in the first six months. The transfer forecasting accuracy is slightly better compared to the Dutch case. This is clear from the baseline RMSEs, which are 0.108 and 0.116 compared to the median baseline of 0.141 in the Netherlands. This is most likely due to the larger scale of the PV installation, as well as the diverse angles of the panels. This causes the output to be smoother compared to smaller installations with one specific angle, making the forecasting exercise easier.
A major difference between BE and NL, however, is that the final month of training lies in the winter for BE and in summer for NL. This would explain why it takes so much longer for a univariate target model in BE to perform decently compared to NL, because it does not have any data from summer months to fine-tune on for the first few months. (5m is August in this case, 6m is July) We will come back to this issue in section 4.4.
@@ 4.3 The effect of weather covariates
Figure 4 shows the results for the Dutch buildings, but now with the addition of three weather variables: relative humidity at 2 meters, direct solar radiation, and diffuse solar radiation. From the plot, we can come to several conclusions. First, the performance increases significantly when including weather variables, the median transfer error decreases from c. 0.117 to 0.085, a performance increase of 27.4%. Second, the performance increase is even more pronounced with target models. When no covariates were included, 7 months of data were needed by the target LSTM to outperform even the naive model on average, but with weather data it only takes 2 months. On average, the transfer model still performs better for the first half year, but the difference is smaller compared to the univariate case. A final observation is that the performance is far more consistent across different sites now, i.e. the confidence band of the target and transfer models is a lot narrower now.
@@ 4.4 The effect of seasonality
To explore this further, we shifted the terminal month in time for different univariate, models. Table 1 shows these results for NL, by focusing specifically on what happens with the RMSE of the target models relative to the daily persistence model. The columns represent the final month of the target data used for training. The earlier in the year we have the terminal month, the less data we have in total, hence the increasing amount of n.a. values. We can see that there is a clear trend: the earlier in the year the final month is, the better. This most likely has to do with the presence of the best performing day inside the training data. Out of the 40 datasets, the month with the highest PV-output hour was in April 21 times and in May eleven times. As soon as this data is available to the model, the performance drastically increases. It is also the case that the main forecasting errors come from underestimating the peaks, which further explains the need for this data to be available during the training phase.
Table 1: Relative RMSE of target model compared to the baseline dependent on the terminal training month for the Dutch case
Dec Nov Oct Sep Aug Jul Jun May
0m 48% 47% 47% 47% 47% 47% 47% 47%
1m 37% 34% 26% 16% 14% 18% 25% 7%
2m 37% 27% 14% 30% 8% 5% 21% 2%
3m 26% 8% -5% -3% 6% -3% 8% -16%
4m 10% -12% -10% 2% 4% 15% -17% -16%
5m 4% -3% -12% 13% 3% -16% -20% n.a.
6m 6% -15% 0% 4% -17% -19% n.a. n.a.
7m -14% -17% -5% -14% -19% n.a. n.a. n.a.
8m -17% -17% -18% -18% n.a. n.a. n.a. n.a.
9m -22% -19% -21% n.a. n.a. n.a. n.a. n.a.
10m -21% -22% n.a. n.a. n.a. n.a. n.a. n.a.
11m -24% n.a. n.a. n.a. n.a. n.a. n.a. n.a.
4.5 The effect of source data mis-specification

### v3

@@ Abstract
Deep learning models have gained increasing prominence in recent years in solar photovoltaic (PV) forecasting. One drawback of these models is that they require a lot of high-quality data to perform well. This is often infeasible in practice, due to poor measurement infrastructure in legacy systems and the rapid build-up of new solar systems across the world. This paper proposes SolNet: a novel, general-purpose, multivariate solar power forecaster, which addresses these challenges by using a two-step forecasting pipeline which incorporates transfer learning from abundant synthetic data generated from PVGIS, before fine-tuning on observational data.
Using actual production data from hundreds of sites in the Netherlands, Australia and Belgium, we show that SolNet improves forecasting performance over data-scarce settings as well as baseline models. We find transfer learning benefits to be the strongest when only limited observational data is available. At the same time we provide several guidelines and considerations for transfer learning practitioners, as our results show that weather data, seasonal patterns and possible misspecification in source location can have a major impact on the results. The SolNet models created in this way are applicable for any land-based solar photovoltaic system across the planet to obtain improved forecasting capabilities.
@@ 1.2 Contributions
Second, we do a thorough deep-dive of various transfer learning scenarios, providing guidance on when and how to use it successfully in solar power generation scenario’s. We look at three performance metrics for better understanding of the value of transfer: initial performance, learning improvement, and asymptotic performance. Figure 1 visualises our hypothesis of the expected performance of the transfer learning. Additionally, we look at (1) the importance of weather data, (2) the impact of the yearly seasonality on transfer learning out-performance, and (3) how misspecification of the source domain can impact the performance of transfer learning.
Figure 1: Complete assessment of transfer performance. The initial performance measures performance of a model without any training in the domain of interest (the target domain). The learning performance indicates the performance gain we can expect over time as more data to train on becomes available. Finally, the asymptotic performance is the ’permanent’ performance differential between the models when sufficient data is available. The shape of the figure is indicative of our hypothesised findings.
@@ 2.1 The forecasting exercise and baseline approach
For this paper we focused on the case of forecasting the next 24 hours, from 00:00 to 23:00 with no gap between inference time and the forecast period. This is a simple reference case to showcase the performance of SolNet. The main reason for choosing this option is that, when using weather forecast data (which we get from the Deutscher Wetterdienst’s Icosahedral Nonhydrostatic EU model (DWD ICON EU) via Open Climate Fix (Open Climate Fix, 2023)), we have to decide on a reference timing. Historical weather forecasts, when and if they are available, typically offer the option to choose from four different reference time slots: 00:00, 06:00, 12:00 and 18:00. We limited ourselves to the 00:00 data, as otherwise the dimensionality of the data would have become too large to work with [^1] Note that DWD forecasts are publicly available several hours after the reference time of 00:00, meaning that a user should either use 2 day-ahead forecasts or a different source for real-life application. We stick to DWD because it is one of the few public sources of historcal weather forecasts..
@@ 2.2 LSTM as a forecast model
$h_{t-1}$ : is the hidden state at time t1t-1
$c_{t-1}$ : is the cell state at time t1t-1
@@ 2.3 Transfer learning
With the deep learning architecture in place, it is possible to explore the effect of transfer learning on solar forecasting in a principled manner. More concretely, the LSTM can be pre-trained on source data, after which it can be fine-tuned using target data to achieve potential improvements over the target model. Transfer learning, formally defined as improving the learning of a target predictive function on the learning task, $\mathcal{F}_{T}(\cdot)$ in the target domain, $\mathcal{D}_{T}$, using knowledge of $\mathcal{D}_{S}$ and $\mathcal{T}_{S}$ (the source domain and the source task respectively) (Pan & Yang, 2010). While $\mathcal{D}_{T}\neq\mathcal{D}_{S}$ and $\mathcal{T}_{T}\neq\mathcal{T}_{S}$, we do assume that, the more similar the two are, the easier the transfer learning task will be. In our case, we keep the source domain as closely related to the target domain as possible by mimicking the location and the PV set-up of our target domain data, in our source domain, but we also provide results for what happens when this condition does not hold (see section 4.4). This is in addition to the inherent misalignment between simulated PVGIS data and observed data.
@@
    Finetuning. The source LSTM can be fine-tuned to the target domain. We keep the architecture intact, and run several additional training iterations using the (limited) target domain data. We can avoid model overfitting on the limited target data by lowering the learning rate, thereby making sure that the model only makes slight adjustments to the previously trained weights. Another popular method to fine-tune deep neural networks is by “freezing” the parameters of certain units or entire layers (see e.g. Zhou et al. (2020b)). In this case, the optimiser can no longer update the parameters in these layers during the optimisation process, but is free to update the remaining parameters.
@@ 2.4 Model evaluation
Performance is measured by way of the root mean squared error (RMSE), in accordance with (Yang et al., 2020), since it penalizes large errors more heavily. This is important to correctly predict the timing and size of the daily peak in output. Another reason to focus on the RMSE is its ubiquity in solar forecasting literature, which makes it easier to compare across studies. In appendix we also include the mean absolute error (MAE) and mean bias error (MBE) for each of the experiments, to provide a more complete picture of model performance. The MAE focuses more on smaller errors compared to the RMSE and the MBE gives insights towards the bias of the models.
In all cases, a full year of target data is kept aside for final evaluation. 80% of the remaining data is used for training and 20% is used to validate the model performance and pick the best performing iteration of the model. The complete workflow of the experiments is presented in figure 2. The first column shows the workflow for a source trained model. The middle column shows the process for a target trained model, skipping the source step and training directly on the target domain instead of finetuning with a lower learning rate and frozen layers. The final column shows the naive seasonal model, which takes the values of the day before as the predictions, requiring no training.
@@ 3.1.1 Residential buildings in the Netherlands
For all 40 locations, the full year of 2020 is available as well as 2021 up until August. We also have a large part of 2019 available for most locations, but unfortunately our weather forecast data is limited to 2020 and 2021, so we cannot go back further for target locations if we want include weather variables. As mentioned, a full year of data is kept aside for evaluation of the final models. Which will thus consist of a part of 2021 and a part of 2020. For the univariate analysis, this leaves us with a full year (for most locations) to evaluate performance, but for the weather analysis we only have January 2020 up until August 2020. This leaves little data for training and validation in the target domain, but this fits well with the core premise of the research that there is often only a (very) limited amount of data available in the target domain before forecasts need to be made.
@@ 3.2 Source domain
For this research we use data generated by PVGIS from 2005 to 2018, mimicking the parameters (latitude, longitude, tilt, azimuth, peak power) of each of the previously discussed solar PV systems (in case these parameters are available) in order to maximize the similarity between source and target domain. So, for example, an actual PV system in EnergyVille, Genk, Belgium with latitude 50.99 and longitude 5.54 a tilt of 33°, point directly south with a peak power of 2.5 kWp would get these exact parameters in the PVGIS tool to generate data from 2005 until 2018.
This data is subsequently used to train the source models we use for transfer learning. These models are then fine-tuned using the sparse observational target data. Even though PVGIS provides data for 2020, we limit ourselves to 2018 while creating source models as otherwise this causes data leakage. Data leakage refers to including knowledge about the future (or data which would realistically not be available) in a model. In this case, data leakage would occur if we were to use the years 2019 and 2020 in the creation of our model and, simultaneously, evaluate the model on 2019 and 2020. Similarly, for the Australian PV systems, for which we have data from 2011 to 2012, we limit the source data until 2010.
@@ 3.3 Weather data
To mitigate this issue, we treat the source and target domain differently. For training the source model, we utilize readily available observed historical weather data. For the target models, we only utilize actual forecast weather data. While this introduces a misalignment between source and target models, it does not cause any data leakage issues and also avoids placing large and potentially expensive forecast availability requirements on SolNet users, or anyone looking to use transfer-learning for their solar power forecasts.
@@ 3.4 Data pre-processing
$x_{\text{max}}=\max(x_{\text{max}},(\text{peak-power}\cdot 0.86))$
@@ 4 Results
In this section, we present results of applying transfer learning to residential PV systems in the Netherlands and Australia, as well as the commercial PV systems in Belgium using univariate data. Next, using the Dutch data, we take a closer look at the impact of including weather covariates. Then we look at the importance of the time of the year when we start gathering target data on model performance. To do this, we change the training period to assess the seasonal impact. Finally, we also assess the impact of data misspecification in the source domain, by increasing the (real) distance between the source and target domain.
4.1 Univariate forecasting of residential and commercial PV systems
The results for univariate forecasting of 40 residential PV systems in the Netherlands are summarised in figure 3(a) (i.e. without incorporating weather observations or forecasts in either the source or target domain). On the x-axis we plot the performance based on the months of data available in the target domain. ”0m” is the case with no target data (cf. 1. Likewise, ”1m” means that one month of target domain data is used for training/finetuning. On the y-axis we plot the scaled RMSE for each of these cases (i.e. the RMSE on the data which has been scaled using the min-max scaler). The thick lines are the overall median [^3] The median values are shown instead of mean values because not all datasets have the same observation period, which greatly impacts the mean estimate. values and the shaded areas represent the range of the individual cases. In other words, the shaded area represents the results of all 40 PV systems, from lowest to highest RMSE. The error of the baseline model (a seasonal naive model taking the values of the 24 hours of the previous day to be the new prediction), is a straight line, which does not improve over time, as it is independent of the amount of data available. Finally, identical figures for the MAE and MBE can be found in figures 6 in the appendix.
@@ 4.1 Univariate forecasting of residential and commercial PV systems
Figure 3: Forecast error as a function of target data availability and use of transfer learning for univariate models. The thick line represents the median performance at each time step. The shaded areas represent the performance of all individual systems, with the outer edges being the worst/best performing systems. The Results indicate transfer learning models consistently outperform both naive baselines and target models (i.e. trained without transfer).
@@
From this plot, we can make four observations. First, the transfer models strictly outperform the baseline models. Second, the target model only starts to outperform the simple baseline once a sufficient amount of data (approx. 4 to 6 months of data depending on the test site) is available to train the model. This is very likely driven by the fact that this time period is sufficient for the model to have seen months with different weather conditions allowing it to generalize. Third, we can see that the asymptotic performance improvement is very small, but persistent, after c. 7 months. Fourth, the transfer model almost always outperforms the target model, until more than a half year of data is available.
@@
Figure 3(b) plots the corresponding results for Australia. Given the larger size of the dataset (299 samples), the shaded area shows the 0.05 to 0.95 percentiles for each month to filter out some of the noise in the outliers. In a similar fashion to the Dutch results, most of the value lies in the initial months. As before, it takes over 7 months for the target model to consistently outperform the naive model here. The learning improvement performance for the transfer model is again small. The median RMSE is 0.135 for the model with no fine-tuning (0m), improving by about 7.4% to 0.125 after 12 months. The target models outperform the transfer models in 8.6% of cases, mostly in the final four months. Interestingly, the asymptotic out-performance is (almost) non-existent for Australia. This could be due to the fact that the source domain is a lot more limited, including only 2005 up until 2010, compared to 2005 up until 2018 for NL.
@@
Figure 3(c) plots the results for the two commercial buildings in Belgium, showing that results are in line with the residential results. As before, the transfer model outperforms the target model, with most of the performance difference in the first six months. The transfer forecasting accuracy is slightly better compared to the Dutch case. This is clear from the baseline RMSEs, which are 0.108 and 0.116 compared to the median baseline of 0.141 in the Netherlands. This is most likely due to the larger scale of the PV installation, as well as the diverse angles of the panels. This causes the output to be smoother compared to smaller installations with one specific angle.
Finally, it’s worth discussing the MAE and MBE of NL, AUS and BE as these metrics tell a more nuanced story than the RMSE. The figures for these metrics can be found in the appendix (6). We can see that the MAE of the target and transfer models is worse, relative to the benchmark, than the RMSE. For NL and BE the MAE of the transfer model is on par, and for AUS it is worse than the benchmark. It is worth noting that the target and transfer models are trained based on the MSE criterion, thus favoring a better RMSE. A good MAE is therefore a byproduct of model training, not the main goal. Still, ideally a transfer learned model outperforms the benchmark both on RMSE and MAE. This does become the case if we include weather data, which we discuss in the next section. In terms of the MBE, we can see that target models have a strong positive bias until sufficient data becomes available (3 to 5 months of data), the transfer models however do not show a strong bias.
@@ 4.2 The effect of weather covariates
Figure 4 shows the results for the Dutch buildings, but now with the addition of three weather variables: relative humidity at 2 meters, direct solar radiation, and diffuse solar radiation. From the plot, we can come to several conclusions. First, the performance increases significantly when including weather variables, the median transfer error decreases from c. 0.117 to 0.085, a performance increase of 27.4%. Second, the performance increase is even more pronounced with target models. When no covariates were included, 7 months of data were needed by the target LSTM to outperform even the naive model on average, but with weather data it only takes 2 months. On average, the transfer model still performs better for the first half year, but the difference is smaller compared to the univariate case, and the asymptotic out-performance is also more limited. A final observation is that the performance is far more consistent across different sites now, i.e. the confidence band of the target and transfer models is a lot narrower now.
@@
When we look at the MAE and MBE of the models that include weather data (see figures 7), the story also changes significantly compared to the models without weather data. When weather data is included as a predictor, we can see that the MAE is in line with the RMSE. The transfer models consistently outperform the benchmark on the MAE, and the target model does so after sufficient data is available to train on (after c. 3 to 5 months). The univariate transfer models thus seem to be adequate in finding the overall shape of the solar curve, limiting the large errors, while the weather data adds in the small errors that happen hour-to-hour. Finally, we can see that the bias in the target models is also gone much faster when weather data is included, while it has no real effect on the bias in transfer models (which was already low).
@@ 4.3 The effect of seasonality
To explore this further, we shifted the terminal month in time for different, univariate models. Table 1 shows these results for NL, by focusing specifically on what happens with the RMSE of the target models relative to the daily persistence model (giving an RMSE skill-score). The columns represent the final month of the target data used for training. The earlier in the year 2019 we have the terminal month, the less data we have in total, hence the increasing amount of n.a. values. We can see that there is a clear trend: the earlier in the year the final month is, the better, up until a point. This most likely has to do with the presence of the best performing day inside the training data. Out of the 40 datasets, the month with the highest PV-output hour was in April 21 times and in May eleven times. As soon as this data is available to the model, the performance drastically increases. It is also the case that the main forecasting errors come from underestimating the peaks, which further explains the need for this data to be available during the training phase.
Table 1: RMSE of target model for NL relative to the baseline dependent on the terminal training month (the columns) and the amount of target data to train and validate on (the rows)
Final Month
Jan/20 Dec/19 Nov/19 Oct/19 Sep/19 Aug/19 Jul/19 Jun/19 May/19 Apr/19 Mar/19 Feb/19
0m 48% 47% 47% 47% 47% 47% 47% 47% 47% 47% 47% 47%
1m 62% 58% 53% 31% 7% 5% 13% 31% 6% 7% 14% 20%
2m 58% 52% 35% -3% 0% 5% 6% 20% 1% -8% 0% n.a.
3m 58% 23% 0% -9% 2% 10% 7% 8% -8% -9% n.a. n.a.
4m 22% -8% -10% -8% 5% 13% 5% 4% -11% n.a. n.a. n.a.
5m -7% -11% -11% -1% 9% 7% -4% -11% n.a. n.a. n.a. n.a.
6m -13% -12% -11% 6% 6% -4% -9% n.a. n.a. n.a. n.a. n.a.
7m -14% -12% -10% 6% -2% -9% n.a. n.a. n.a. n.a. n.a. n.a.
8m -13% -11% -12% 1% -7% n.a. n.a. n.a. n.a. n.a. n.a. n.a.
9m -13% -9% -4% -7% n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
10m -14% -9% -8% n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
11m -13% -11% n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
12m -11% n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a. n.a.
4.4 The effect of source data misspecification